Two particles are fixed on an x axis. Particle 1 of charge 65.5 μC is located at x = -22.0 cm; particle 2 of charge Q is located at x = 31.6 cm. Particle 3 of charge magnitude 24.6 μC is released from rest on the y axis at y = 22.0 cm. What is the value of Q if the initial acceleration of particle 3 is in the positive direction of (a) the x axis and (b) the y axis?
Answer #1

As we know that
Start with diagram.
Angle formed between pt1 and pt 3, A = tan-1(22/22) = 45
Angle formed between pt2 and pt 3, B = tan-1(22/31.6) = 34.8457
Distance from pt1 to pt3, R = sqrt(22^2 + 22^2) = sqrt(968)

(Note, you only need R^2 later)
Distance from pt2 to pt3, S = sqrt(22^2 + 31.6^2) = 38.50402576

(Note, you only need S^2 later)

Call the charges Q1, Q and Q3.

If the initial acceleration of particle 3 is along the positive x-direction, then the y components of F13 and F23 must cancel.
(Q1.Q3/R^2)sinA = (Q.Q3/S^2)sinB
Q3 cancels. You know all the values except Q. Simplify and calculate Q.

Q = 1.003178512*10^-4(sinA/sinB)

Q = 1.241500919*10^-4 C